OpenStudy (anonymous):

solve the inequality (2x+15)/x < x

OpenStudy (anonymous):

where did u get x^3?

jhonyy9 (jhonyy9):

2x+15x<x*2 x*2 -17x >0 x(x-17) >0

OpenStudy (anonymous):

2x+15<x^2 <=> x^2-2x-15>0 D=4+60=64 x1=6 x2=-2

OpenStudy (anonymous):

I am not sure,sorry If I wrong.

OpenStudy (anonymous):

thx for all ur help y'alls

OpenStudy (llort):

my mistake, I typed it in wrong. 15/x+2<x

OpenStudy (anonymous):

Do you want me yo continue or you are okay?

OpenStudy (llort):

here is your answer...

jhonyy9 (jhonyy9):

when x>17

OpenStudy (anonymous):

so i did (2x+15) < x^2

OpenStudy (anonymous):

@jhonyy9 where did you find the 15x ?

OpenStudy (sgadi):

Hi Alexix, The procedure is right. But there is a small mistake in the first step. The inequality, gives two possible answers based on the sign of x. \[(2x+15)/x<x\] The work given by you is correct for x>0. But for x<0, you get another solution which is \[(2x+15)/x<x\Rightarrow 2x+15>x^2\]